# LookAhead: Preventing DeFi Attacks via Unveiling Adversarial Contracts

### v4

@@ 5.2. Pruned Semantic-Control Flow Tokenization (PSCFT)
Based on state-of-the-art EVM contract binary lifter and analysis framework Gigahorse (Grech et al., 2019), we design a pipeline for 1) Constructing intra-contract Control Flow Graphs (CFG) and external call chains based on Intermediate Representation (IR), 2) Aggregating and extracting dataset features from EVM bytecode, and 3) Generating textual summaries for the smart contracts in our dataset.
@@ 5.2.1. Implementation Feature Extraction
To enable the extraction of implementation features (cf. §4.2.2) simultaneously with the generation of the IR, we implemented a custom-designed Souffle Datalog program based on existing rules from Gigahorse. First, by utilizing the Statement_Opcode rule, which maps each statement to its corresponding opcode, we can directly extract the number of occurrences of critical opcodes, such as DELEGATECALL and SELFDESTRUCT, which corresponds to features $delegate\_call\_count$ and $selfdestruct\_count$. Next, we leverage the FunctionEntry rule, which contains all function entry points (identifiers), with the help of PublicFunctionSelector, which maps public functions to their selectors, to determine the total number of public and private functions. This combination also helps identify flashloan callback functions. Finally, to analyze the external function calls within a contract, we design a rule called ExternalCallResolved. This rule provides details about all resolved external calls made within the contract, including their selectors and signatures. Additionally, we implemented logic to handle special cases, such as identifying fallback functions and transfers of native tokens, based on the arguments and values passed to the calls. With the help of ExternalCallResolved, we can keep track of token-related calls and their related information including their total count, maximum occurrences, and average usage.
@@ 5.2.2. CFG Construction and Pruning
Based on the IR of a contract, we construct a function-by-function visualization $F=\{F^{private}\cup F^{public}\}$ of the internal structure of contract methods, each consisted of control flow blocks and programming statements. Functions have 4-byte selector derived from the Keccak hash (SHA3) of their canonical names and can be either private or public. Each function contains multiple basic blocks (BB) chained together ($\{\mathcal{F}_{i}=\{BB^{i}_{0},BB^{i}_{1},\cdots,BB^{i}_{|\mathcal{F}_{i}|}\}\}$), and each block contains multiple statements. Blocks and statements are identifiable via unique numerical IDs, for example, $90x11b$ 9. The relationship between two blocks can be discovered by following the predecessor and successor IDs associated with each block, each block can have multiple predecessors and successors, hence we can create a directed graph $G_{i}=(B_{i},A_{i})$ such that the set of nodes $B=\{\mathcal{BB}^{i}_{j}:\mathcal{BB}^{i}_{j}\in\mathcal{F}_{i}\}$ where $\mathcal{F}_{i}\in F$, and $(\mathcal{BB}^{i}_{x},\mathcal{BB}^{i}_{y})$ is an arc in $A_{i}$ if $\mathcal{BB}^{i}_{x}$ is a predecessor of $\mathcal{BB}^{i}_{y}$ and $\mathcal{BB}^{i}_{y}$ is a successor of $\mathcal{BB}^{i}_{x}$. Statements could be constant/variable declarations, internal/external function calls, function terminations, memory copies, storage initialization, etc., we focus on constructing a CFG based on function call statements.
The pruning process starts with the sorting and reordering of functions based on function types (public first, then private) and their canonical names. Gigahorse can only recover the names for some of the public functions and external call methods, hence if the derivation of canonical names from function selectors failed, we will leave them as-is and recover their names in the semantic recovery stage. For any private function $\mathcal{F}^{private}_{j}\in F^{private}$, we give it a canonical name of $InternalFunction_{i}$. Next, we remove all statements not related to function calls from every block $\mathcal{BB}^{i}_{j}\in B_{i}$. Then, we perform a depth-first search (DFS) over $G$ to remove any block $\mathcal{BB}^{i}_{k}$ that contains no statements (i.e. $|\mathcal{BB}^{i}_{k}|=0$) from $B_{i}$, along with its incoming and outgoing arcs from $A_{i}$, and update the the predecessor and successor IDs for its neighbours. Finally, we assign canonical names $BB^{i}_{j},j\in[0,|B|-1]$ to basic blocks that are still left in $B_{i}$.
@@ 5.2.3. Semantic Recovery
For an external call statement like the one shown in section PSCFT and Feature Construction in Figure 6, we aim to recover: 1) An address label for 0xBA12222222228d8Ba445958a
75a0704d566BF2C8
, 2) A human-readable name for the target method 0x5c38449e.
Similar to our fund source labelling algorithm, we attempt to match the target address with an address associated with a known label in Etherscan’s address label cloud and public name tags. For contracts that cannot be matched with an existing label, we check if they are open source, when they are, we use their contract name as the label. Contract developers might choose to store the target address in the EVM storage or as a variable/constant, in which case we use the same technique as (Kong et al., 2023) to run a Datalog program that finds the slot location storing the address value and retrieves the value from the target storage. If this process did not result in a matching label, we assign the address a default label UnknownTarget.
Function selectors are generated by hashing the function signatures with a secure hashing algorithm, and the hashing process is a one-way trap door. However, thanks to the deterministic nature of hash functions, third-party service such 4byte (4by, 2024) are able to construct and maintain a large database of publicly known hash-to-name mappings. Therefore, in hopes to include as much function names and as little unknown functions in the PSCFT as possible, we extend Gigahorse’s name recovery ability with 4byte’s service by sending requests to their APIs upon failed recovery by Gigahorse. If the recovery still fails, we fallback to the default label UnknownFunc.
If the aforementioned process was completed successfully, the external call statement will be converted to Balancer.flashLoan(…args). And we can produce a textual summary for smart contracts containing control-flow and semantic call information simply by concatenating $\mathcal{F}_{i}=\{\mathcal{BB}^{i}_{0},\cdots,\mathcal{BB}^{i}_{|\mathcal{F}_{i}-1|}\}\in F$ together in the form shown in Figure 6.3, Semantic Recovery.
@@ 5.3.1. Semantic Transformer
To ensure the model’s generalization ability and improve training efficiency, we implement regularization techniques such as early-stopping and dropouts to prevent model over-fitting. Overall, this transformer model provides a semantic view of DeFi smart contracts.
@@ 5.3.2. Ensemble Classifier
Aiming to embed the features selected based on our empirical observations into the semantic view of the transformer model, we design an ensemble classifier that merges the classifier models trained on basic features with the transformer model trained on PSCFT together to produce a meta classifier model with better predictive performance.
Candidate Classifiers. Once all features have been extracted, we can proceed to train a supervised learning classifier using a variety of ML algorithms, outputting multiple candidate models for further evaluation. These models include classical approaches such as Logistic Regression $\mathcal{M}^{classifier}_{LR}$ and Decision Trees $\mathcal{M}^{classifier}_{DT}$, ensemble learning methods such as bagging-based Random Forest $\mathcal{M}^{classifier}_{RF}$ and XGBoost $\mathcal{M}^{classifier}_{XGBM}$ based on boosting.
@@
For the fund source feature, which consists of categorical data (ref. §4.2.1) with low cardinality, we apply one-hot encoding. To facilitate better comparability among input values in numerical terms of varying dimensions, we standardize the dataset using Z-score normalization. This process brings each feature value around to the centre of 00 with a standard deviation of $1$.
After training has completed for each of the candidate classifiers, we evaluation their performance using metrics such as Recall and F1-Score to determine an optimal model $\mathcal{M}^{classifier}_{optimal}\in\{\mathcal{M}^{classifier}_{LR},\cdots,\mathcal{M}^{classifier}_{XGBM}\}$ for the meta classifier training and inference in the next stage.
Meta Classifier. Using the stacked generalization (stacking) machine learning algorithm, we design a blender that stacks $\mathcal{M}^{transformer}$ and $\mathcal{M}^{classifier}_{optimal}$ together to form a higher level model $\mathcal{M}$ known as a meta classifier. By feeding the predictions produced by the two base models into the meta model, we are able to 1) Combine the strengths of each model and automatically learn the optimal way to construct the inference result based on model significance, 2) Optimize the final classification boundaries to achieve more accurate detection of adversarial contracts.
We use K-Nearest Neighbors (KNN) to as our meta classifier $\mathcal{M}$ due to its simplicity, effectiveness, and ability to make predictions based on the proximity of data points in the feature space. KNN is well-suited for combining the outputs of different classifiers such as the transformer model and one of the candidate classifiers, because it is able to combine the complementary strengths of both models by analyzing their predictions and finding the optimal decision boundary.
@@ 6. Evaluation
We implement LookAhead in Python using open source machine learning frameworks PyTorch and Tensorflow to achieve both CPU and GPU-accelerated training and inference of our classifier models. All experiments are conducted on a server running a 3.70GHz Intel Xeon processor with 32 threads and a NVIDIA A100 GPU with 40GB of VRAM.
@@ 6.1. Experiment Setup
6.1.1. Dataset
Token and proxy contracts are two of the primary types of contracts used on blockchains, while they are commonly used by developers, most adversaries do not rely on those types of contracts for their core attack logic. To identify contract types, we perform a simple check based on PSCFT to determine whether a contract follows known standards such as ERC20 and ERC1967. By analyzing whether the contracts match common standards, we found that in the dataset we built, $184,386$ benign samples were token contracts, $10,037$ were proxy contracts, and none of the adversary contracts belonged to these two categories. Therefore, we remove them from both our training and testing dataset the same way as it was in previous system Forta (For, 2024).
A distinguishing aspect of our research when compared to traditional supervised machine learning problems is the explicit chronological order of contract deployment. Contracts deployed over a different periods of time may exhibit varying characteristics and vulnerabilities, meaning that adversarial contracts may evolve and change over time as well. One of LookAhead’s objectives is to provide a ML classifier model that is pre-trained on past DeFi incidents but capable of detecting potential future attacks. Consequently, traditional cross-validation techniques, which assume input values are independent with respect to each other, are not suitable for our evaluation. Instead, we use the earliest $80\%$ of the contracts for training and reserve the latest $20\%$ for testing. Out of the $80\%$ of contracts used as training data, we take the last $25\%$ of most recent samples as the training set for the meta classifier, and the rest for the transformer and candidate models.
@@ 6.1.2. Research Questions
NLP Classifier Transformer 0.9876 0.7429 0.7027 0.7222 0.0057
Forta - 0.8778 0.5520 0.6208 -
Meta Classifier KNN 0.9953 0.9286 0.8667 0.8966 0.0016
Table 1. Performance breakdown of different models.
@@
    RQ1. How effective is LookAhead’s system of classifiers in detecting the adversarial contracts used in DeFi attacks, especially when compared to previous state-of-the-art solutions?
@@
    RQ2. How efficient can the LookAhead system be in identifying adversarial contracts, can it produce a valid response quick enough so that defense actions could take place?
@@
    RQ3. How practical is it for individuals or organizations to deploy the LookAhead system and what does the performance of LookAhead look like in real-world scenarios?
@@ 6.2.2. Experiment Results
Table 1 presents our experiment results for evaluating the performance of LookAhead’s ML models. We calculate the performance metrics of various models used in the process of forming the LookAhead detection system, including $4$ candidate classifiers, a transformer-based classifier, and a meta classifier. Among the candidate classifiers, XGBoost achieves the highest F1-score ($0.8553$), while Random Forest (RF) has the lowest recall ($0.7432$). Overall, the XGBoost-based classifier outperforms others, demonstrating the best performance across most metrics compared to other candidates. Its false positive rate reaches as low as $0.0041$, and it exhibits a high effective detection rate for adversarial samples. Other models also perform well. Except the one based on Decision Tree (DT), the F1-scores of all other classifiers exceed $0.79$, proving the validity of our method. The meta classifier, K-nearest Neighbors (KNN), combines the optimal candidate model XGBoost with the transformer to achieve the best overall performance with the highest accuracy ($0.9953$), F1-Score ($0.8966$), and the lowest FPR ($0.0016$).
6.2.3. Comparison with the Baseline
There have been studies that analyze smart contracts to detect contract vulnerabilities, but to the best of our knowledge, this is the first academic work that focus on analyzing smart contracts for the detection of adversarial contracts used by attackers across all types of DeFi attacks. The goal of both research directions is to prevent attackers from exploiting vulnerable contracts for profits, the only different is that our method aim to stop attacks from causing damage to potential targets while other works help potential targets reduce the possibility of being attacked. However, we emphasize that existing works on detecting vulnerabilities often focus on specific types of attacks/vulnerabilities, offering limited protection capabilities. Therefore, when comparing our results with the baseline, our goal is to: 1) Highlight the general applicability of LookAhead’s system in detecting all types of DeFi attacks, 2) Demonstrate the performance of LookAhead against previous state-of-art solutions.
@@ 6.2.3. Comparison with the Baseline
Universality. We compare the performance of LookAhead with a previous state-of-the-art academic work (Yang et al., 2024) that specifically detects reentrancy attack contracts to demonstrate the universality of the LookAhead system. When our trained meta classifier model was tested on all the reentrancy related adversarial contracts in our test dataset, we observed that LookAhead raised an alert for $83.33\%$ of the attack contracts, which is very similar to the results presented by (Yang et al., 2024). This observation along with the experiment results presented in §6.2.2 show that despite being trained on contracts used in a variety of attacks of different types (cf. §4.1.1), LookAhead still achieves comparable performance in detecting any specific types of DeFi attacks.
@@
Performance. The only similar work that uses machine learning for adversarial contract detection that was made public so far is the Forta Network (For, 2024). However, their approach relies solely on EVM bytecode while our method is based on a combination of multi-dimensional features and semantic tokenization (PSCFT) of smart contracts. We use Forta as the baseline for comparing the performance of our ML models.
We consider Forta an NLP-based method, hence we perform a direct comparison between their model and our transformer model first, thereafter we measure how much benefits can our meta classifier approach provide over Forta’s implementation. As the results in Table 1 indicate, our transformer alone can perform better and achieve higher F1-Socre than Forta, demonstrating the advantages of our Pruned Semantic-Control Flow Tokenization design.

### v6

@@ 5.2. Pruned Semantic-Control Flow Tokenization (PSCFT)
Based on EVM bytecode binary lifter framework Gigahorse (Grech et al., 2019), we design a pipeline for 1) Constructing intra-contract Control Flow Graphs (CFG) and external call chains based on Intermediate Representation (IR), 2) Aggregating and extracting dataset features from EVM bytecode, and 3) Generating textual summaries for the smart contracts in our dataset.
@@ 5.2.1. Implementation Feature Extraction
To enable the extraction of implementation features (see §4.2.2) simultaneously with the generation of the IR, we implemented a custom-designed Souffle Datalog program based on existing rules from Gigahorse. 1) First, by utilizing the Statement_Opcode rule, which maps each statement to its corresponding opcode, we can directly extract the number of occurrences of critical opcodes, such as DELEGATECALL and SELFDESTRUCT, which corresponds to features $delegate\_call\_count$ and selfdestruct_count. 2) Next, we leverage the FunctionEntry rule, which contains all function entry points (identifiers), with the help of PublicFunctionSelector, which maps public functions to their selectors, to determine the total number of public and private functions. This combination also helps identify flashloan callback functions. 3) Finally, to analyze the external function calls, we design the ExternalCallResolved rule, providing details about all resolved external calls made within the contract, including their selectors and signatures. With the help of ExternalCallResolved, we can keep track of token-related calls and their related information including their total count, maximum occurrences, and average usage.
@@ 5.2.2. CFG Construction and Pruning
Based on the IR of a contract, we construct a function-by-function visualization =$\mathcal{F}=\{\mathcal{F}^{private}\cup\mathcal{F}^{public}\}$ of the internal structure of contract methods, each consisted of control flow blocks and programming statements. Functions have 4-byte selector derived from the Keccak hash (SHA3) of their canonical names and can be either private or public. Each function contains multiple basic blocks (BB) chained together ($\{F_{i}=(BB^{i}_{0},BB^{i}_{1},\cdots,BB^{i}_{|F_{i}|})\}$), and each block contains multiple statements. Blocks and statements are identifiable via unique numerical IDs, for example, $90x11b$ 9. The relationship between two blocks can be discovered by following the predecessor and successor IDs associated with each block, each block can have multiple predecessors and successors, hence we can create a directed graph $G_{i}=(B_{i},A_{i})$ such that the set of nodes $B_{i}=\{BB^{i}_{j}:BB^{i}_{j}\in F_{i}\}$ where $F_{i}\in\mathcal{F}$, and $(BB^{i}_{x},BB^{i}_{y})$ is an arc in $A_{i}$ if $BB^{i}_{x}$ is a predecessor of $BB^{i}_{y}$ and $BB^{i}_{y}$ is a successor of $BB^{i}_{x}$. Statements could be constant/variable declarations, internal/external function calls, function terminations, memory copies, storage initialization, etc., we focus on constructing a CFG based on function call statements.
The pruning process starts with the sorting and reordering of functions based on function types (public first, then private) and their canonical names. Gigahorse can only recover the names for some of the public functions and external call methods, hence if the derivation of canonical names from function selectors failed, we will leave them as-is and recover their names in the semantic recovery stage. For any private function $F^{private}_{j}\in\mathcal{F}^{private}$, we give it a canonical name of $InternalFunction_{i}$. Next, we remove all statements not related to function calls from every block $BB^{i}_{j}\in B_{i}$. Then, we perform a depth-first search (DFS) over $G$ to remove any block $BB^{i}_{k}$ that contains no statements (i.e. $|BB^{i}_{k}|=0$) from $B_{i}$, along with its incoming and outgoing arcs from $A_{i}$, and update the the predecessor and successor IDs for its neighbours. Finally, we assign canonical names $BB^{i}_{j},j\in[0,|B|-1]$ to basic blocks that are still left in $B_{i}$.
@@ 5.2.3. Semantic Recovery
For an external call statement like the one shown in §PSCFT and Feature Construction in Figure 5, we aim to recover: 1) An address label for 0x5777d92f208679db4b9778590
fa3cab3ac9e2168
, 2) A human-readable name for the target method 0x490e6cbc.
Similar to our fund source labelling algorithm, we consider to match the target address with an address associated with a known label in chain explorers’ address label cloud and public name tags. For contracts that cannot be matched with an existing label, we check if they are open source, when they are, we use their contract name as the label. Contract developers might choose to store the target address in the EVM storage or as a variable/constant, in which case we use the same technique as (Kong et al., 2023) to run a Datalog program that finds the slot location storing the address value and retrieves the value from the target storage. If this process did not result in a matching label, we assign the address a default label UnknownTarget.
Function selectors are generated by hashing the function signatures with a secure hashing algorithm, and the hashing process is a one-way trap door. However, thanks to the deterministic nature of hash functions, third-party service such 4byte (4by, 2024) are able to construct and maintain a large database of publicly known hash-to-name mappings. Therefore, in hopes to include as much function names and as little unknown functions in the PSCFT as possible, we attempt to extend Gigahorse’s name recovery ability with 4byte’s service by sending a request to their API upon failed recovery by Gigahorse. If the recovery still fails, we fallback to the default label UnknownFunc.
If the aforementioned process was completed successfully, the external call statement will be converted to UniswapV3.flash(…args). And we can produce a textual summary for smart contracts containing control-flow and semantic call information simply by concatenating $F_{i}=(BB^{i}_{0},\cdots,BB^{i}_{|F_{i}-1|})$ $\in\mathcal{F}$ together in the form shown in Figure 4.
@@ 5.3.1. Transformer
To ensure the model’s generalization ability and improve training efficiency, we implement regularization techniques such as early-stopping and dropouts to prevent model over-fitting. Overall, this transformer model provides a semantic view of smart contracts.
@@ 5.3.2. Ensemble Classifier
Aiming to embed the features selected based on our empirical observations into the semantic view of the transformer model, we design an ensemble classifier that merges the classifier models trained on empirical features with the transformer model trained on PSCFT together to produce a meta classifier model with better predictive performance.
Candidate Classifiers. Once all features have been extracted, we can proceed to train a supervised learning classifier using a variety of ML algorithms, outputting multiple candidate models for further evaluation. These models =$\mathcal{M}=\{M^{classifier}_{LR},M^{classifier}_{DT},M^{classifier}_{RF},M^{classifier}_{XGBM}\}$ include classical methods such as Logistic Regression and Decision Trees, ensemble learning methods such as bagging-based Random Forest and XGBoost which is based on boosting.
@@
For the fund source feature, which consists of categorical data (see §4.2.1) with low cardinality, we apply one-hot encoding. To facilitate better comparability among input values in numerical terms of varying dimensions, we standardize the dataset using Z-score normalization. This process brings each feature value around to the centre of 00 with a standard deviation of $1$.
After training has completed for each of the candidate classifiers, we evaluate their performance using a function $f:\mathcal{M}\rightarrow\mathbb{R}$ that incorporates metrics such as Recall and F1-Score to determine an optimal model $M^{classifier}$ such that $M^{classifier}=\text{argmax}_{M\in\mathcal{M}}f(M)$ for the meta classifier training and inference in the next stage.
Meta Classifier. Using the stacked generalization (stacking) machine learning algorithm, we design a blender that stacks $M^{transformer}$ and $M^{classifier}$ together to form a higher level model $M^{meta}$ known as a meta classifier. By feeding the predictions produced by the two base models into the meta model, we are able to 1) Combine the strengths of each model and automatically learn the optimal way to construct the inference result based on model significance, 2) Optimize the final classification boundaries to achieve more accurate detection of adversarial contracts.
Given that the meta classifier operates on low-dimensional inputs (only the two probability values between 00-$1$ outputted from the $M^{classifier}$ and $M^{transformer}$, indicating the likelihood of a contract being adversarial) and without incorporating any additional information about the contracts, employing a simple model is both sufficient and effective. This not only ensures efficiency but also mitigates the risk of overfitting by limiting model complexity. Accordingly, we only consider simple models, including K-Nearest Neighbors (KNN), Logistic Regression (LR), Support Vector Machines (SVM), and Decision Trees (DT), and evaluate their performance to select the most suitable meta classifier in experiments.
@@ 6. Evaluation
We implement LookAhead in Python using open source machine learning frameworks PyTorch and Tensorflow to achieve both CPU and GPU-accelerated training and inference of our classifier models. All experiments are conducted on a server running a 3.70GHz Intel(R) Xeon(R) processor with 32 threads and a NVIDIA A100 GPU with 40GB of VRAM.
@@ 6.1. Experiment Setup
6.1.1. Datasets
Token and proxy contracts are two of the primary types of contracts used on blockchains, while they are commonly used by developers, most adversaries do not rely on those types of contracts for their core attack logic. To identify contract types, we perform a simple check based on PSCFT to determine whether a contract follows known standards such as ERC20 and ERC1967. By analyzing whether the contracts match common standards, we found that in the dataset we built, $184,386$ benign samples were token contracts, $10,037$ were proxy contracts, and none of the adversary contracts belonged to these two categories. Therefore, we remove them from both our training and testing dataset the same way as it was in previous work Forta (For, 2024).
A distinguishing aspect of our research when compared to traditional supervised machine learning problems is the explicit chronological order of contract deployment. Contracts deployed over a different periods of time may exhibit varying characteristics and vulnerabilities, meaning that adversarial contracts may evolve and change over time as well. One of LookAhead’s objectives is to provide a ML classifier model that is pre-trained on past DeFi incidents but capable of detecting potential future attacks. Consequently, we sort the contract samples chronologically and use the earliest $80\%$ of the contracts for training and reserve the latest $20\%$ for testing. Out of the $80\%$ of contracts used as training data, we take the last $25\%$ of most recent samples as the training set for the meta classifier, and the rest for the transformer and candidate models.
@@ 6.1.2. Research Questions
NLP Classifier Forta - 0.8778 0.5520 0.6208 -
Transformer 0.9876 0.7429 0.7027 0.7222 0.0057
Meta Classifier LR 0.9950 0.9683 0.8133 0.8841 0.0006
DT 0.9913 0.7901 0.8533 0.8205 0.0054
SVM 0.9947 0.9143 0.8533 0.8828 0.0019
KNN 0.9953 0.9286 0.8667 0.8966 0.0016
@@
    RQ1. How effective is LookAhead in detecting the adversarial contracts used in DeFi attacks?
@@
    RQ2. How do different empirical features contribute to model decision-making and how does PSCFT improve performance compared to using raw IR?
@@
    RQ3. How efficient can LookAhead be in identifying adversarial contracts?
  4. 4)
    RQ4. How practical is LookAhead  in real-world scenarios?
@@ 6.2.2. Evaluation Results
Table 1 presents our experimental results evaluating the performance of various ML models (§5.3) for the LookAhead system, including four candidate classifiers, a transformer-based classifier, and four meta classifiers. Among the candidate classifiers, XGBoost achieves the highest F1-score ($0.8553$), while Random Forest has the lowest recall ($0.7432$). Overall, the XGBoost-based classifier outperforms other candidate classifiers, demonstrating the best performance across most metrics compared to others. It achieves a false positive rate as low as $0.0041$ while maintaining a high effective detection rate for adversarial samples. Other models also perform well. Except the one based on DT, the F1-scores of all other classifiers exceed $0.79$, proving the validity of our method. Building upon the results of the candidate models, different meta classifiers were constructed by combining the optimal candidate model XGBoost with the transformer. The results show that three of the four meta classifiers achieved higher F1-scores compared to the standalone XGBoost classifier. Among them, the KNN-based classifier demonstrated the best performance, achieving the highest F1-score ($0.8966$) while maintaining a low FPR ($0.0016$).
6.2.3. Comparison with Previous Works
We compare our approach with two state-of-the-art baselines: Forta (For, 2024), a NLP-based method for detecting all adversarial contracts, and BlockWatchDog (Yang et al., 2024), a specialized work focused exclusively on detecting reentrancy attack contracts.
@@ 6.2.3. Comparison with Previous Works
Universality. To evaluate the universality of the LookAhead system, we compare its performance against BlockWatchDog, which specifically targets reentrancy attack contracts. We test our meta classifier on all the reentrancy related adversarial contracts in our test dataset, we observe that LookAhead successfully identified $83.33\%$ of them, which is on par with the results presented by (Yang et al., 2024). This observation along with the experiment results presented in §6.2.2 show that LookAhead not only generalizes well across diverse types of DeFi attacks but also maintains competitive performance when applied to specific attack categories.
@@
Performance. The only similar work that uses machine learning for adversarial contract detection publicly available so far is the Forta Network (For, 2024). We use Forta as the baseline for comparing the performance of our ML models.